1- A ball with mass M, moving horizontally at 2.8 m/s, collides elastically with a block with mass 3.6M that is initially hanging at rest from the ceiling on the end of a 58-cm wire. Find the maximum angle through which the block swings after it is hit, in degrees.

2- A 0.15 kg glider is moving to the right on a frictionless, horizontal air track with a speed of 0.72 m/s. It has a head-on collision with a 0.30 kg glider that is moving to the left with a speed of 0.36m/s. Suppose the collision is elastic. Find the magnitude of the final velocity (in m/s) of the 0.150kg glider.
Answer #1

1.

mb = mass of the ball = M

mB = mass of the block = 3.6 M

vbo = initial velocity of ball before collision = 2.8 m/s

vb = velocity of ball after collision

VB = velocity of the block after collision

Using conservation of momentum

mb vbo = mb vb + mB VB

M (2.8) = M vb + (3.6) M VB

2.8 = vb + (3.6) VB

vb = 2.8 - (3.6) VB Eq-1

using conservation of kinetic energy

mb v2bo = mb v2b + mB V2B

M (2.8)2 = M v2b + (3.6) M V2B

(2.8)2 = v2b + (3.6) V2B

Using eq-1

(2.8)2 = (2.8 - (3.6) VB )2 + (3.6) V2B

VB = 1.22 m/s

h = height gained by the block

Using conservation of energy for the block after the collision

Kinetic energy at the bottom just after collision = potential energy gained by the block

(0.5) mB VB2 = mB g h

h = (0.5) VB2 /g

h = (0.5) (1.22)2 /9.8

h = 0.076 m

= angle through which the block swings

L = length of the wire = 58 cm = 0.58 m

height gained is also given as

h = L (1 - cos)

0.076 = 0.58 (1 - cos)

cos = 0.87

= Cos-1(0.87)

= 29.5 deg